What is 1/3 of 6?

Found 2 solutions by usyim88hk, sschoentrup:
Answer by usyim88hk(157)

It is 6 times 1 and then divide by 3
which is (6*1)/3 = 6/3 = 2

Answer by sschoentrup(1)


1/3 x 6/1 = 6/3 = 2/1 = 2